OpenStudy (anonymous):

I just want to check my answer to this problem: 30-8x=5(x+6) does x= 0/26?

OpenStudy (anonymous):

If 0 is divided by something isn't it always equal to 0?

OpenStudy (anonymous):

0 is correct

OpenStudy (anonymous):

30-8x=5x+30 0=5x+8x 13x=0 x=0

OpenStudy (anonymous):

sorry not 300 its just 30